OpenStudy (anonymous):

Find the value of the determinant. |2 7 -3| |-1 5 -4| |6 9 0 |

OpenStudy (fifciol):



OpenStudy (jdoe0001):

$$ \left[ \begin{matrix} 2 & 7 & -3\\ -1 & 5& -4\\ 6 & 9 & 0 \end{matrix} \right] \implies -3(+1) \left[ \begin{matrix} -1& 5\\ 6 & 9 \end{matrix} \right] -4(-1) \left[ \begin{matrix} 2& 7\\ 6& 9 \end{matrix} \right]\\ -3(-9-30)+4(18-42)\\ -3(-39)+4(24) \implies 21 $$

OpenStudy (jdoe0001):

$$ \left[ \begin{matrix} 2 & 7 & -3\\ -1 & 5& -4\\ 6 & 9 & 0 \end{matrix} \right] \implies -3(+1) \left[ \begin{matrix} -1& 5\\ 6 & 9 \end{matrix} \right] -4(-1) \left[ \begin{matrix} 2& 7\\ 6& 9 \end{matrix} \right]\\ -3(-9-30)+4(18-42)\\ -3(-39)+4(-24) \implies 21 $$

OpenStudy (jdoe0001):

better, tis -24